OpenStudy (anonymous):

-Shannon's scooter has a piece of tape stuck to the tire. If the tire has a diameter of 18 inches, how far does the piece of tape travel in 72° of rotation? 2π 9π 2 pi over 5 18 pi over 5

OpenStudy (campbell_st):

well the circumference of a circle in terms of diameter is \[C = \pi \times d\] the tape will travel the circumference of the tire... so find the circumference

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

confused O.o

OpenStudy (campbell_st):

ok so the circumference of the tire is \[C = 18\pi\] what fraction of a circle is 72 degrees \[\frac{72}{360}\] so the distance traveled is the circumference x the fraction... \[\frac{72}{360} \times 18\pi\] you'll need to simplify the answer

OpenStudy (anonymous):

the answer is 18π over 5